OpenStudy (anonymous):

Short Answer: Type your answer in the space below each question. Show your work when applicable. 6. The velocity of sound in air is given by the equation =20√(273+t) , where v is the velocity in meters per second and t is the temperature in degrees Celsius. Find the temperature when the velocity is 329 meters per second by graphing the equation. Round the answer to the nearest degree

OpenStudy (goformit100):

Is it a Physics Question

OpenStudy (anonymous):

Nooo its not.. its algebra 2

OpenStudy (anonymous):

Algebra eh?

OpenStudy (mathmale):

Yes, the problem does stem from Physics, showing up in an Algebra course. But everything necessary to know is contained in the presentation of this problem. v(t)=20√(273+t).

OpenStudy (mathmale):

Supposing that you are told that the velocity at one point in time is 329 m/sec; your job is to determine the time at which that is true.

OpenStudy (mathmale):

Good. How are you going to solve that for time, T?

OpenStudy (mathmale):

Although the problem asks you to find T graphically, you could certainly solve this problem algebraically and compare the result (T = ?) to the solution you obtain through graphing.

OpenStudy (anonymous):

329=20sqrt(273+T) 329/20=sqrt(273+T) 16.45=sqrt(273+T) 16.45^2=273+T

OpenStudy (anonymous):

now im lost

OpenStudy (mathmale):

so far, so good! Get out your calculator again. Type in 16.45 and then whichever key will square that 16.45 for you. Write down your result, and then equate it to 273 + T.

OpenStudy (anonymous):

16.45^2=270.6

OpenStudy (mathmale):

Yes, and so you have the equation 270.6 = 273 + T. Please solve for T. To accomplish this, subtract 273 from both sides. Note: This will give you a T that is negative. Is that OK or not OK? Why?

OpenStudy (anonymous):

ty mathmale i had the same question

OpenStudy (anonymous):

-2.4 and no because the negatives change to a positive

OpenStudy (mathmale):

Glad to be of help. @LissaBaby989: I agree with your result, T (actually, t) = - 2.4. But think: what are we measuring with T or t? Can that quantity be negative?

OpenStudy (anonymous):

A negative number taken to an even power gives a positive result

OpenStudy (mathmale):

True, but we have already squared both sides, before attempting to solve for t. So that reason does not apply here.

OpenStudy (anonymous):

im confused how the last steps go.. :/

OpenStudy (anonymous):

-2.4-273=T =-275.4?

OpenStudy (mathmale):

Hold. Reviewing the original problem, I saw that T represents temperature, not time. So there's no difficulty if T comes out to be negative. -2.4 degrees Celcius is roughly 25 degrees F, which is certainly possible. Find that temperature, T, to the nearest degree. If any part of the last few posts here is unclear for you, point them out, so that we can discuss them.

OpenStudy (anonymous):

so T =25 degress F?

OpenStudy (mathmale):

Yes. I need to ask you whether or not that sounds reasonable to you. Does it?

OpenStudy (anonymous):

Not sure.